Determine the temperature, volume, internal energy, enthalpy, entropy and density of 2kg of
steam at a pressure of 2MM/m2 and with a dryness fraction of 0.85.
1
Expert's answer
2019-04-25T09:31:04-0400

For this problem we need h,S diagram


On this diagram find a point where across two lines: pressure 20 bar and dryness fraction 0,85

For it we have

Entropy per kg

"S=5.7565 (\frac{kJ}{kg\cdot K})"

Enthalpy per kg

"i=2516 (\frac{kJ}{kg})"


Volume per kg


"\frac{V}{m}=0.0878 (\frac{m^3}{kg})"


Find internal energy from this equation

"i=U+pV"

Internal energy

"U=i-pV"

For temperature use this equation

"T=\frac{di}{dS}""T=\frac{2550-2516}{5.7565-5.826}-273=216 ^\circ C"

Density

"\rho=\frac{m}{V}=11.4 (\frac{kg}{m^3})"

So, for 2 kg mass of steam and pressure 20 bar, we have

Volume

"V=0.0878*2=0.1756 (m^3)"

Internal energy

"U=5032-2\cdot 10^{3} \cdot 0.1756=4681 (kJ)"

Enthalpy

"i=2516*2=5032 (kJ)"

Entropy

"S=5.7565*2=11.513 (\frac{kJ}{K})"


Density

"\rho=11.4 (\frac{kg}{m^3})"